OpenStudy (anonymous):

Determine the vertices of a triangle whose sides have the mid-points (1,0), (2, -1), and (3,1). Thanks in advance. :)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

do we get any choices?

OpenStudy (amistre64):

i draw a triangle; but I cant see a way to anchor the lines

OpenStudy (amistre64):

well, maybe...

OpenStudy (amistre64):

any ideas to help my memory out with :)

OpenStudy (anonymous):

I'm also having confusion on finding the points.. :))

OpenStudy (amistre64):

well; point are defined as (go over this many, go up this many) (go over to x=# , go to y=#)

OpenStudy (anonymous):

I see...

OpenStudy (amistre64):

heres the point

OpenStudy (anonymous):

mark the points x1,y1 x2, y2 and x3, y3 now (x1+x2)/2 = 1 (x2+x3)/2= 2 (x3+x1)/2=3

OpenStudy (anonymous):

uve got three equations and three variables, solve for obtaining x1, x2, x3

OpenStudy (anonymous):

now do it the same way for the y-coordinates :)

OpenStudy (anonymous):

Ohh alright.. :)

OpenStudy (anonymous):

hope this does it...

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

yeah, that helps my focus :)

OpenStudy (anonymous):

Cool, awesome! Thanks to both of you. :))

OpenStudy (anonymous):

thank you mr mistretta

OpenStudy (anonymous):

oh no problem bonny miss

OpenStudy (anonymous):

:)))

OpenStudy (anonymous):

parentheses brought to life by gorgeous smiles...unprecedented experience...:)

OpenStudy (anonymous):

loool :D